# Heterogeneous Treatment Effects and

### v3

@@ Appendix E Comparison to Mediation Analysis
In this section, we compare our framework connecting HTEs and mechanisms to mediation analysis. It is important to note that the two frameworks are built on different principles and objects. The main purpose of mediation analysis is to identify and estimate various average causal mediation effects (ACMEs). Identification of these effects relies on the assumption of sequential ignorability (Imai and Yamamoto, 2013). On the other hand, our framework aims to infer the activation of a mechanism (for at least one unit in the sample) by using heterogeneous treatment effects, which instead, relies on exclusion assumptions that we propose (Assumptions 2-3). The following DAGs facilitate our discussion of the differences in these approaches.
@@
Consider first the case with multiple independent causal mechanisms in Figure A.3. In the left DAG, treatment $Z$ indirectly affects outcome $Y$ through two channels $M_{1}$ and $M_{2}$, and may also directly affect $Y$. To non-parametrically identify average indirect effect mediated by $M_{1}$, the key part of the sequential ignorability is $Y_{i}(z^{\prime},m_{1},M_{2i}(z^{\prime}))\perp M_{1i}|Z_{i}=z$. The assumption is challenging to interpret and cannot be guaranteed by most experimental designs because it involves cross-index independence, from $z$ to $z^{\prime}$. Graphically, sequential ignorability requires that all variables $U$ should be observed and included in the adjustment set. Another important implication of the assumption is that $U$ cannot be affected by the treatment $Z$, i.e., the dashed line is not allowed. When these assumptions hold, mediator $M_{1}$ must be measured in order to estimate the ACME.
@@
There are many other differences between the two methods. For example, in the mediation analysis, mediators must be measurable and measured while these measurements are not required by our framework. However, in our framework, researchers must have a measured candidate MDV $X$ that is believed to satisfy Assumptions 2-3. Also, when using HTEs to detect mechanisms, researchers need to pay more attention to whether $Y$ is directly affected outcome. Even though we have emphasized their differences, two frameworks also have shared features. For example, both require that the total causal effect of $Z$ on $Y$ is identified. This is explicitly assumed by sequential ignorability $\{Y_{i}(z,m_{1i},m_{2i}),M_{1i},M_{2i}\}\perp Z_{i}$ and implicitly assumed in our framework.
@@ Appendix F Proofs of Propositions
We prove propositions under general multi-mechanism cases, as stated in the Appendix C. All propositions in the main text are all re-stated in the special case of one mechanism. Assumption 3 is only invoked in the multiple mechanism generalization.
@@ F.1 Proof of Proposition 1
Suppose that $Y$ is directly affected by mechanism $j$ and Assumptions 2 and 3 hold with respect to $X_{k}$. If HTEs exist with respect to $X_{k}$, then $X_{k}\in\textbf{X}_{j}^{MDV}$ for mechanism $j$.
@@
$CATE_{L(Y)}(X_{k}=x)$ $=E_{X_{\neg k}}[L(Y)|Z=z,X_{k}=x]-E_{X_{\neg k}}[L(Y)|Z=z^{\prime},X_{k}=x]$ (35)
$=E_{X_{\neg k}}[L(DE(z,z^{\prime};X_{k}=x)+\sum_{j=1}^{J}IE_{j}(z,z^{\prime};X_{k}=x))]$ (36)
$=L\{E_{X_{\neg k}}[DE(z,z^{\prime};X_{k}=x)+\sum_{j=1}^{J}IE_{j}(z,z^{\prime};X_{k}=x)]\}$ (37)
$=L[ADE(z,z^{\prime};X_{k}=x)+\sum_{j=1}^{J}AIE_{j}(z,z^{\prime};X_{k}=x)]$ (38)
Equation (36) follows from the linearity of expectations and the decomposition of the total effect in 11. Equation (37) is guaranteed by the linearity of $L$.
Then, under exclusion assumptions 2 and 3, we can express:
$CATE_{L(Y)}(X_{k}=x)-CATE_{L(Y)}(X_{j}=x^{\prime})=L[AIE_{j}(z,z^{\prime};X_{k}=x)-AIE_{j}(z,z^{\prime};X_{k}=x^{\prime})]$ (39)
@@ F.2 Proof of Propostions 2
Suppose that $Y$ is directly affected by mechanism $j$ and Assumptions 1 and 2 hold. If no HTEs exist with respect to $X_{k}$, at least one of the following must be true:
@@
Proof by contrapositive. Suppose not, which means $\textbf{X}_{j}^{MDV}$ is non-empty and for some $x,x^{\prime}\in\mathbb{R}$, $IE_{j}(X_{k}=x)\neq IE_{j}(X_{k}=x^{\prime})$. Then:
$L\{E_{X_{\neg k}}[IE_{j}(z,z^{\prime};X_{k}=x)\}$ $\neq L\{E_{X_{\neg k}}[IE_{j}(z,z^{\prime};X_{k}=x^{\prime})]\}$ (40)
We then can reconstruct $CATE_{L(Y)}(X_{k}=x)\neq CATE_{L(Y)}(X_{k}=x^{\prime})$ from (40):
$L\{E_{X_{\neg k}}[IE_{j}(z,z^{\prime};X_{k}=x)]\}$ $\neq L\{E_{X_{\neg k}}[IE_{j}(z,z^{\prime};X_{k}=x^{\prime})]\}$ (41)
$E_{X_{\neg k}}[L(DE(X_{k}=x)+IE_{j}(X_{k}=x)+IE_{i\neq j}(X_{k}=x)]$ $\neq E_{X_{\neg k}}[L(DE(X_{k}=x^{\prime})+\sum_{j=1}^{J}IE_{j}(z,z^{\prime};X_{k}=x)]$ (42)
$CATE_{L(Y)}(X_{j}=x)$ $\neq CATE_{L(Y)}(X_{j}=x^{\prime})$ (43)
Note $E_{X_{\neg k}}[DE(X_{k}=x^{\prime})]=ADE(X_{k}=x)$ and $E_{X_{\neg k}}[IE_{j}(z,z^{\prime};X_{k}=x)]=AIE_{j}(X_{k}=x^{\prime})$. Equation (42) follows because $ADE(X_{k}=x)=ADE(X_{k}=x^{\prime})$ and $AIE_{i\neq j}(X_{k}=x)=AIE_{i\neq j}(X_{k}=x^{\prime})$ under exclusion assumptions 2 and 3. We find HTEs with respect to $X_{k}$.
@@ F.3 Proof of Proposition 3
Suppose that observed outcome $h(Y)$ is a non-linear mapping of directly-affected outcome $Y$ and Assumptions 1 and 2 hold. If HTEs exist with respect to $X_{k}$, then $X_{k}\in\textbf{X}^{R}.$
@@
For additional intuition about how non-linear transformations of $Y$ affect HTE, we use the following Lemma.
Lemma 1.
Given pre-treatment variables {$_{1},X_{2},...,X_{n}\}=\textbf{X}$ X and outcome variable $Y$. Variables {$_{1},...,X_{m}\}\subset\textbf{X}$ X are MDVs for mechanism j (denote them as the set $\textbf{X}_{j}^{MDV}$ and the remaining as the set $\textbf{X}_{j}^{non-MDV}$) if and only if there exits function $g_{1}(\cdot)$ and non-additively separable function $g_{2}(\cdot)$, and $Y$ satisfies:
@@
A function, $F(X_{1},X_{2})$, will be called additively separable if it can written as $f_{1}(X_{1})+f_{2}(X_{2})$ for some functions $f_{1}(X_{1})$ and $f_{2}(X_{2})$. Note further that:
@@
    Equation (48) in the above theorem should be understood as
@@
    The non-additively separable function $g_{2}(\textbf{X}_{j}^{MDV},Z)$ can take the form $g_{3}(T)+g_{4}(\textbf{X}_{j}^{MDV},T)$ for some function $g_{3}(\cdot)$ and non-additively separable function $g_{4}(\cdot)$.
@@
$CATE(X=x)-CATE(X=x^{\prime})=E[h_{2}(\textbf{X}_{j}^{MDV},Z)|X=x]-E[h_{2}(\textbf{X}_{j}^{MDV},Z^{\prime})|X=x^{\prime}]$ (49)
@@ F.4 Extension of Proposition 4
Suppose that observed outcome $h(Y)$ is a non-linear mapping of directly-affected outcome $Y$ and Assumptions 1 and 2 hold. If HTEs do not exist with respect to $X_{k}$, then $_{k}\in\textbf{X}.$ X .
Proposition 8 indicates that if there exist no HTEs for the indirectly-affected outcome, $X_{k}$ can be any relevant or non-relevant covariate. Now we provide a stronger version of Proposition 8 by imposing assumptions about the directly-affected outcome, $Y$, and the form of the non-linear transformation $h(Y)$. These assumptions permit additional learning from the lack of HTEs in this case.
In practice, most indirectly-affected outcomes are discrete variables, such as voting behavior, survey responses, or choices. Let us consider the following non-linear transformation of the directed affected outcome $Y$:
$h(Y)=\begin{cases}y_{1}&Y\in(-\infty,c_{1}]\\ y_{2}&Y\in(c_{1},c_{2}]\\ ...\\ y_{q}&Y\in(c_{q-1},\infty)\\ \end{cases}$ (50)
Here, will assume $y_{i}\in\mathbb{R}$ in equation (50) has no substantive interpretation. In practice, values of $y_{i}$ are typically normalizations, that are arbitrarily determined by the researcher. As such, the value is independent of model parameters.
@@
$_{k}\in\textbf{X}$ X is effective if $\exists i\in\{1,2,...,q\}$ and $x,x^{\prime}\in X_{k}$ such that $p_{i}(x;z,z^{\prime})\neq p_{i}(x^{\prime};z,z^{\prime})$.
@@
Appendix G Strengthening Assumptions
In the main text, Propositions 3 and 4 indicate that for indirectly-affected outcomes, the existence or non-existence of HTEs are not generally informative about mechanism activation. In this section, we explore the conditions under which invoking stronger assumptions can provide more information.
Recall the basic problem with indirectly-affected outcomes. In Figure A.5, $Y$ is the directly-affected outcome, and $h(Y)$ is the non-linear transformation of the directly-affected outcome. Our main result shows that $X^{k+1}$ can also induce HTEs even though $X^{k+1}$ does not detect the mechanism.

### v4

@@ Appendix E Comparison to Mediation Analysis
In this section, we compare our framework connecting HTEs and mechanisms to mediation analysis. It is important to note that the two frameworks are built on different principles and objects. The main purpose of mediation analysis is to identify and estimate various average causal mediation effects (ACMEs). Identification of these effects relies on the assumption of sequential ignorability (Imai and Yamamoto, 2013). On the other hand, our framework aims to infer the activation of a mechanism (for at least one unit in the sample) by using heterogeneous treatment effects, which instead, relies on exclusion assumptions that we propose (Assumptions 1-2). The following DAGs facilitate our discussion of the differences in these approaches.
@@
Consider first the case with multiple independent causal mechanisms in Figure A.6. In the left DAG, treatment $Z$ indirectly affects outcome $Y$ through two channels $M_{1}$ and $M_{2}$, and may also directly affect $Y$. To non-parametrically identify average indirect effect mediated by $M_{1}$, the key part of the sequential ignorability is $Y_{i}(z^{\prime},m_{1},M_{i2}(z^{\prime}))\perp M_{i1}|Z_{i}=z$. The assumption is challenging to interpret and cannot be guaranteed by most experimental designs because it involves cross-index independence, from $z$ to $z^{\prime}$. Graphically, sequential ignorability requires that all variables $U$ should be observed and included in the adjustment set. Another important implication of the assumption is that $U$ cannot be affected by the treatment $Z$, i.e., the dashed line is not allowed. When these assumptions hold, mediator $M_{1}$ must be measured in order to estimate the ACME.
@@
There are many other differences between the two methods. For example, in the mediation analysis, mediators must be measurable and measured while these measurements are not required by our framework. However, in our framework, researchers must have a measured candidate MDV $X$ that is believed to satisfy Assumptions 1-2. Also, when using HTEs to detect mechanisms, researchers need to pay more attention to whether $Y$ is directly affected outcome. Even though we have emphasized their differences, two frameworks also have shared features. For example, both require that the total causal effect of $Z$ on $Y$ is identified. This is explicitly assumed by sequential ignorability $\{Y_{i}(z,m_{i1},m_{i2}),M_{i1},M_{i2}\}\perp Z_{i}$ and implicitly assumed in our framework.
@@ Appendix F Proofs of Propositions
We prove propositions under general multi-mechanism cases, as stated in the D.
@@ F.1 Proof of Proposition 1
Assumptions 1-2 are sufficient and generically necessary for $CATE(X_{k}=x)-CATE(X_{k}=x^{\prime})=AIE^{Y}_{1}(z,z^{\prime};X_{k}=x)-AIE^{Y}_{1}(z,z^{\prime};X_{k}=x^{\prime})$.
@@
$CATE^{L(Y)}(X_{k}=x)$ $=E_{X_{\neg k}}[L(Y)|Z=z,X_{k}=x]-E_{X_{\neg k}}[L(Y)|Z=z^{\prime},X_{k}=x]$ (23)
$=E_{X_{\neg k}}[L(DE^{Y}(z,z^{\prime};X_{k}=x)+\sum_{j=1}^{J}IE^{Y}_{j}(z,z^{\prime};X_{k}=x))]$ (24)
$=L\{E_{X_{\neg k}}[DE^{Y}(z,z^{\prime};X_{k}=x)+\sum_{j=1}^{J}IE^{Y}_{j}(z,z^{\prime};X_{k}=x)]\}$ (25)
$=L[ADE^{Y}(z,z^{\prime};X_{k}=x)+\sum_{j=1}^{J}AIE^{Y}_{j}(z,z^{\prime};X_{k}=x)]$ (26)
Equation (24) follows from the linearity of expectations and the decomposition of the total effect in 19. Equation (25) is guaranteed by the linearity of $L$.
Therefore,
$CATE^{L(Y)}(X_{k}=x)-CATE^{L(Y)}(X_{k}=x^{\prime})$ (27)
$=$ $L[ADE^{Y}(z,z^{\prime};X_{k}=x)-ADE^{Y}(z,z^{\prime};X_{k}=x^{\prime})]$ (28)
$+$ $L[\sum_{j=2}^{J}AIE^{Y}_{j}(z,z^{\prime};X_{k}=x)-AIE^{Y}_{j}(z,z^{\prime};X_{k}=x^{\prime})]$ (29)
$+$ $L[AIE^{Y}_{1}(z,z^{\prime};X_{k}=x)-AIE^{Y}_{1}(z,z^{\prime};X_{k}=x^{\prime})]$ (30)
For sufficiency, under exclusion assumptions 1 and 2, equation (28) and (29) are zero.
[…]
Case (1) is equivalent to assumptions 1-2. This can be realized if covariate $X_{k}$ does not ‘modify’ the average direct and indirect effects.
For case (2), covariate $X_{k}$ does modify the average direct or/and indirect effects. Generically, it can take any real numbers. However, the probability that two real numbers exactly offset, $\Delta ADE^{Y}=-\Delta AIE^{Y}$ is 0.
Therefore, we conclude that generically, case (1), i.e., assumptions 1-2, hold.
F.2 Proof of Proposition 2
Proposition 7.
Suppose Assumptions 1-2 hold with respect to $X_{k}$ for outcome $Y$. If HTEs exist with respect to $X_{k}$, then $X_{k}\in\textbf{X}^{MDV}$ for mechanism $M_{j}$.
Proof.
Follow the same proof in the Proposition 6, we have $CATE^{L(Y)}(X_{k}=x)=L[ADE^{Y}(z,z^{\prime};X_{k}=x)+\sum_{j=1}^{J}AIE^{Y}_{j}(z,z^{\prime};X_{k}=x)]$.
Then, under exclusion assumptions 1 and 2, we can express:
$CATE^{L(Y)}(X_{k}=x)-CATE^{L(Y)}(X_{j}=x^{\prime})=L[AIE^{Y}_{j}(z,z^{\prime};X_{k}=x)-AIE^{Y}_{j}(z,z^{\prime};X_{k}=x^{\prime})]$ (31)
@@ F.3 Proof of Proposition 3
Suppose Assumptions 1-2 hold with respect to $X_{k}$ for outcome $Y$. If no HTEs exist with respect to $X_{k}$, at least one of the following must be true: If no HTEs exist with respect to $X_{k}$, at least one of the following must be true:
@@
Proof by contrapositive. Suppose not, which means $\textbf{X}_{j}^{MDV}$ is non-empty and for some $x,x^{\prime}\in\mathbb{R}$, $AIE^{Y}_{j}(X_{k}=x)\neq AIE^{Y}_{j}(X_{k}=x^{\prime})$. Then:
We then can reconstruct $CATE^{L(Y)}(X_{k}=x)\neq CATE^{L(Y)}(X_{k}=x^{\prime})$.
$L\{AIE^{Y}_{j}(z,z^{\prime};X_{k}=x)\}$ $\neq L\{AIE^{Y}_{j}(z,z^{\prime};X_{k}=x^{\prime})\}$ (32)
$L\{ADE^{Y}(X_{k}=x)+AIE^{Y}_{j}(X_{k}=x)+AIE^{Y}_{i\neq j}(X_{k}=x)\}$ $\neq L\{ADE^{Y}(X_{k}=x^{\prime})+\sum_{j=1}^{J}AIE^{Y}_{j}(z,z^{\prime};X_{k}=x)\}$ (33)
$CATE^{L(Y)}(X_{j}=x)$ $\neq CATE^{L(Y)}(X_{j}=x^{\prime})$ (34)
Equation (33) follows because $ADE^{Y}(X_{k}=x)=ADE^{Y}(X_{k}=x^{\prime})$ and $AIE^{Y}_{i\neq j}(X_{k}=x)=AIE^{Y}_{i\neq j}(X_{k}=x^{\prime})$ under exclusion assumptions 1 and 2. We find HTEs with respect to $X_{k}$.
@@ F.4 Proof of Proposition 4
Suppose Assumptions 1-2 hold with respect to $X_{k}$ for outcome $Y$. Let observed outcome $h(Y)$ be a transformed outcome relative to $Y$. If HTEs exist with respect to $X_{k}$ for outcome $h(Y)$, then $X_{k}\in\textbf{X}^{R}.$
@@
The preceding proof does not explicitly invoke the definition of $h(\cdot)$ or rely on the assumptions concerning $Y$. A more direct – though considerably more tedious – proof can be carried out by introducing $\delta$ and performing the required integrations.
A more important question is whether $X_{k}\in\mathbf{X}^{R}\setminus\mathbf{X}^{MDV}$, meaning that $X_{k}$ belongs to $\mathbf{X}^{R}$ but not to $\mathbf{X}^{MDV}$. We can demonstrate that this is possible with a simple example. Suppose that we are interested in how a mobilization treatment, $Z_{i}\in\{0,1\}$, affects citizens’ decision to vote. Consider two covariates that predict turnout: $X_{1}\sim\mathcal{N}(0,1)$ and $X_{2}\in\{0,1\}$. We will further assume that $X_{1}$ is a MDV for the unique mechanism, such that:
$M(Z;\textbf{X})$ $=(1+Z)X_{1}$
Potential voters’ utility from voting is given by:
$U(Z,X)$ $=M(Z;X_{1})+X_{2}=(1+Z)X_{1}+X_{2}$
$h(U(Z,X))$ $=\begin{cases}1&\text{ if }(1+Z)X_{1}+X_{2}\geq 0\\ 0&\text{ else }\end{cases}$
In this case, Assumptions 1-2 hold by construction since there is only one mechanism. Now, suppose that a researcher mistakenly thought that $X_{2}$ was a MDV for a mechanism (either $M$ or an inert mechanism). Since we have constructed the data generating process, we know that it is not: $X_{2}\in\textbf{X}^{R}\setminus\textbf{X}^{MDV}$. Evaluating the CATE of $Z$ on turnout when $X_{2}=1$, we have:
$CATE(X_{2}=1)$ $=E[h(U(Z=1,X))-h(U(Z=0,X))|X_{2}=1]$
$=\Pr(2X_{1}+1>0)-\Pr(X_{1}+1>0)$
$=\Phi(-1)-\Phi(-\frac{1}{2})$
$\approx-0.15$
For additional intuition about how non-linear transformations of $Y$ affect HTE, consider the following decomposition: Given pre-treatment variables $\{X_{1},X_{2},...,X_{n}\}=\textbf{X}$. Let variables $\{X_{1},...,X_{m}\}\subset\textbf{X}$ are MDVs for mechanism j (denote them as the set $\textbf{X}_{j}^{MDV}$ and the remaining as the set $\textbf{X}_{j}^{non-MDV}$).
@@
where $g_{1}(\cdot)$ is any function while $g_{2}(\cdot)$ is non-additively separable. A function, $F(X_{1},X_{2})$, will be called additively separable if it can written as $f_{1}(X_{1})+f_{2}(X_{2})$ for some functions $f_{1}(X_{1})$ and $f_{2}(X_{2})$.
Note that:
@@
    (39) should be understood as
@@
    The non-additively separable function $g_{2}(\textbf{X}_{j}^{MDV},Z)$ can take the form $g_{3}(Z)+g_{4}(\textbf{X}_{j}^{MDV},Z)$ for some function $g_{3}(\cdot)$ and non-additively separable function $g_{4}(\cdot)$.
@@
$CATE(X=x)-CATE(X=x^{\prime})=E[h_{2}(g_{2}(\textbf{X}_{j}^{MDV},Z))|X=x]-E[h_{2}(g_{2}(\textbf{X}_{j}^{MDV},Z^{\prime}))|X=x^{\prime}]$ (40)
@@ F.5 Extension of Proposition 5
Suppose Assumptions 1-2 hold with respect to $X_{k}$ for outcome $Y$. Let observed outcome $h(Y)$ be a transformed outcome relative to $Y$. If HTEs do not exist with respect to $X_{k}$ for outcome $h(Y)$, then $X_{k}\in\textbf{X}.$
Proof.
The result follows simply because if $X_{k}\notin\textbf{X}$, then $X_{k}=\emptyset$. ∎
Proposition 10 indicates that if there exist no HTEs for the transfomed outcome, $X_{k}$ can be any relevant or non-relevant covariate. The proof again does not explicitly invoke the definition of $h(\cdot)$ or rely on the assumptions concerning $Y$. Simple examples can show that $X_{k}$ can belong to $\textbf{X}^{MDV}$ and $\textbf{X}^{non-MDV}$. Now we provide a stronger version of Proposition 10 by imposing assumptions about the outcome, $Y$, and the form of the non-linear transformation $h(Y)$. These assumptions permit additional learning from the lack of HTEs in this case.
In practice, most transformed outcomes are discrete variables, such as voting behavior, survey responses, or choices. Let us consider the following non-linear transformation of the directed affected outcome $Y$:
[…]
Here, will assume $y_{i}\in\mathbb{R}$ in (41) has no substantive interpretation. In practice, values of $y_{i}$ are typically normalizations, that are arbitrarily determined by the researcher. As such, the value is independent of model parameters.
@@
$X_{k}\in\textbf{X}$ is effective if $\exists i\in\{1,2,...,q\}$ and $x,x^{\prime}\in X_{k}$ such that $p_{i}(x;z,z^{\prime})\neq p_{i}(x^{\prime};z,z^{\prime})$.
@@
Appendix G Examples from the literature
G.1 Description of study context, design, measures
Table A.2 provides a summary of the context, design, and measures used in the empirical studies detailed in Section 4. Note that the Arias et al. (2019) does not focus on heterogeneous treatment effects by partisan affiliation, though partisan affiliation (bias) is a parameter in their model.
Eggers (2014) Arias et al. (2019) Anduiza, Gallego, and Muñoz (2013) de Figueiredo, Hidalgo, and Kasahara (2023)
Context UK parliamentary elections in 2010 Mexican municpal elections in 2015 Spain 2010 (survey) Sao Paulo, Brazil mayoral runoff in 2008
Design Observational study Field experiment Survey experiment Field, survey experiments
Treatment Binary indicator of sufficient implication in 2009 expenses scandal coded from media reports Dissemination of information on municipal use of intergovernmental infrastructure transfer in municipalities with high or low malfeasant spending Newspaper article-style vignette about corrupt mayor that randomizes partisan affiliation. Flyer distribution about one of two candidates in runoff having convictions involving impropriety in office (field experiment). Same information distributed to untreated voters after election (survey experiment).
Outcome(s) Incumbent vote share (constituency) and British Election Survey self-reported vote for incumbent Precinct-level incumbent party vote share. Perceived seriousness of corruption allegation on 1-10 scale. Precinct-level voting outcomes for each candidate (field experiment). Rating of support for each candidate on 0-10 scale (survey experiment).
Partisanship Indicator for Labour-Conservative constituency. (Not measured directly) Respondent closeness/affiliation to party in vignette. Pre-treatment support for runoff candidate
Table A.2: Design of the four studies discussed in Section 4. Note that some studies include additional outcome variables or different moderators (instead of or in addition to partisanship).
G.2 Justification of theoretical representations
The four examples that we use in Section 4 develop and communicate their theories in different ways. We attempt to remain as faithful as possible to the underlying theories when translating these theoretical assumptions into the models we present in Figure 3. We describe these “translations” below (with documentation from the original studies):
  1. 1.
    Eggers (2014): Presents a formal framework to illustrate the argument. In the framework, a voter ($i$) will vote for the incumbent from party $p_{I}$ over challenger from party $p_{C}$ if:
    $u^{i}(p_{I})-u^{i}(p_{C})\geq c_{I}^{(i)},$ Equation (1), p. 445
    where $u^{i}(\cdot)$ is the utility of having a politician of a given party in power and $c_{I}$ is the incumbent “candidate’s perceived corruption level” (p. 445). The individual-level superscript on $c_{I}^{(i)}$ is our addition.
    Translating to our notation, $a_{i}\equiv u^{i}(p_{I})-u^{i}(p_{C})$ measures alignment with the incumbent party. Suppose further that $c_{I}^{i}$ may at the individual level as a function of individual exposure to information and an aversion to corruption. While we use the parameter $c_{i}$ slightly differently, in this case: